\begin{thebibliography}{10}

\bibitem{agent2d}
Akiyama, H.:
\newblock {Agent2D Base Code}.
\newblock \url{http://www.rctools.sourceforge.jp} (2010)

\bibitem{POWH12}
Prokopenko, M., Obst, O., Wang, P., Held, J.:
\newblock {Gliders2012: Tactics with action-dependent evaluation functions}.
\newblock In: RoboCup 2012 Symposium and Competitions: Team Description Papers,
  Mexico City, Mexico, June 2012. (2012)

\bibitem{gliders2013tdp}
Prokopenko, M., Obst, O., Wang, P., Budden, D., Cliff, O.:
\newblock {Gliders2013: Tactical analysis with information dynamics}.
\newblock In: RoboCup 2013 Symposium and Competitions: Team Description Papers,
  Eindhoven, The Netherlands, June 2013. (2013)

\bibitem{Manual2002}
Chen, M., Dorer, K., Foroughi, E., Heintz, F., Huang, Z., Kapetanakis, S.,
  Kostiadis, K., Kummeneje, J., Murray, J., Noda, I., Obst, O., Riley, P.,
  Steffens, T., Wang, Y., Yin, X.:
\newblock Users Manual: {R}obo{C}up Soccer Server --- for Soccer Server Version
  7.07 and Later.
\newblock The {R}obo{C}up Federation. (February 2003)

\bibitem{moore}
Moore, E., Obst, O., Prokopenko, M., Wang, P., Held, J.:
\newblock Gliders2012: Development and competition results.
\newblock CoRR \textbf{abs/1211.3882} (2012)

\bibitem{Noda2003}
Noda, I., Stone, P.:
\newblock {The RoboCup Soccer Server and CMUnited Clients: Implemented
  Infrastructure for MAS Research}.
\newblock Autonomous Agents and Multi-Agent Systems \textbf{7}(1--2)
  (July--September 2003)  101--120

\bibitem{balch}
Balch, T.:
\newblock Hierarchic social entropy: An information theoretic measure of robot
  group diversity.
\newblock Autonomous Robots \textbf{8}(3) (June 2000)  209--238

\bibitem{IAAI2000}
Stone, P., Riley, P., Veloso, M.:
\newblock Defining and using ideal teammate and opponent models.
\newblock In: Proceedings of the Twelfth Annual Conference on Innovative
  Applications of Artificial Intelligence. (2000)

\bibitem{Butler2001}
Butler, M., Prokopenko, M., Howard, T.:
\newblock Flexible synchronisation within {RoboCup} environment: A comparative
  analysis.
\newblock In: RoboCup 2000: Robot Soccer World Cup IV, London, UK, Springer
  (2001)  119--128

\bibitem{Reis2001}
Reis, L.P., Lau, N., Oliveira, E.:
\newblock Situation based strategic positioning for coordinating a team of
  homogeneous agents.
\newblock In: Balancing Reactivity and Social Deliberation in Multi-Agent
  Systems, From RoboCup to Real-World Applications (selected papers from the
  ECAI 2000 Workshop and additional contributions), London, UK, Springer (2001)
   175--197

\bibitem{Prokopenko02}
Prokopenko, M., Wang, P.:
\newblock Relating the entropy of joint beliefs to multi-agent coordination.
\newblock In Kaminka, G.A., Lima, P.U., Rojas, R., eds.: RoboCup 2002: Robot
  Soccer World Cup VI. Volume 2752 of Lecture Notes in Computer Science.,
  Springer (2003)  367--374

\bibitem{Prokopenko03}
Prokopenko, M., Wang, P.:
\newblock Evaluating team performance at the edge of chaos.
\newblock In Polani, D., Browning, B., Bonarini, A., Yoshida, K., eds.: RoboCup
  2003: Robot Soccer World Cup VII. Volume 3020 of Lecture Notes in Computer
  Science., Springer (2003)  89--101

\bibitem{cliff2013towards}
Cliff, O., Lizier, J., Wang, R., Wang, P., Obst, O., Prokopenko, M.:
\newblock Towards quantifying interaction networks in a football match.
\newblock In: RoboCup 2013: Robot Soccer World Cup XVII, Springer (2013)

\bibitem{PLOW11}
Prokopenko, M., Lizier, J.T., Obst, O., Wang, X.R.:
\newblock Relating {Fisher} information to order parameters.
\newblock Physical Review E \textbf{84}(4) (2011)  041116

\bibitem{mac03}
MacKay, D.J.:
\newblock Information Theory, Inference, and Learning Algorithms.
\newblock Cambridge University Press, Cambridge (2003)

\bibitem{ForemanPW03}
Foreman, M., Prokopenko, M., Wang, P.:
\newblock Phase transitions in self-organising sensor networks.
\newblock In Banzhaf, W., Christaller, T., Dittrich, P., Kim, J.T., Ziegler,
  J., eds.: ECAL. Volume 2801 of Lecture Notes in Computer Science., Springer
  (2003)  781--791

\bibitem{liz08c}
Lizier, J.T., Prokopenko, M., Zomaya, A.Y.:
\newblock The information dynamics of phase transitions in random {B}oolean
  networks.
\newblock In Bullock, S., Noble, J., Watson, R., Bedau, M.A., eds.: Proceedings
  of the Eleventh International Conference on the Simulation and Synthesis of
  Living Systems (ALife XI), Winchester, UK, Cambridge, MA, MIT Press (2008)
  374--381

\bibitem{pro09}
Prokopenko, M., Boschetti, F., Ryan, A.J.:
\newblock An information-theoretic primer on complexity, self-organization, and
  emergence.
\newblock Complexity \textbf{15}(1) (2009)  11--28

\bibitem{Prokopenko13}
Prokopenko, M.:
\newblock Information dynamics at the edge of chaos: Measures, examples, and
  principles.
\newblock In: ALIFE, IEEE (2013)  140--144

\bibitem{liz08a}
Lizier, J.T., Prokopenko, M., Zomaya, A.Y.:
\newblock Local information transfer as a spatiotemporal filter for complex
  systems.
\newblock Physical Review E \textbf{77}(2) (2008)  026110

\bibitem{liz14a}
Lizier, J., Prokopenko, M., Zomaya, A.:
\newblock {A Framework for the Local Information Dynamics of Distributed
  Computation in Complex Systems}.
\newblock In Prokopenko, M., ed.: Guided Self-Organization: Inception. Volume~9
  of Emergence, Complexity and Computation.
\newblock Springer Berlin Heidelberg (2014)  115--158

\bibitem{liz11b}
Lizier, J.T., Pritam, S., Prokopenko, M.:
\newblock Information dynamics in small-world {B}oolean networks.
\newblock Artificial Life \textbf{17}(4) (2011)  293--314

\bibitem{liz11b2}
Lizier, J.T., Pritam, S., Prokopenko, M.:
\newblock Computational capabilities of small-world boolean networks.
\newblock In: Advances in Artificial Life, ECAL 2011, Proceedings of The
  Eleventh European Conference on the Synthesis and Simulation of Living
  Systems (ECAL 2011), Paris, 2011, MIT Press (2011)  463--464

\bibitem{schr00}
Schreiber, T.:
\newblock Measuring information transfer.
\newblock Physical Review Letters \textbf{85}(2) (2000)  461--464

\end{thebibliography}

\documentclass[runningheads,a4paper]{llncs}

\usepackage{amssymb,amsmath}

\usepackage{arydshln}

\usepackage{graphicx}

\usepackage[breaklinks,colorlinks=true,citecolor=blue]{hyperref}

\DeclareMathOperator*{\argmax}{arg\,max}

\begin{document}

\mainmatter  % start of an individual contribution

% first the title is needed
\title{``Infographics'' team: Selecting Control Parameters via Maximal Fisher Information}

% a short form should be given in case it is too long for the running head
%\titlerunning{Gliders2014}

% the name(s) of the author(s) follow(s) next
%
% NB: Chinese authors should write their first names(s) in front of
% their surnames. This ensures that the names appear correctly in
% the running heads and the author index.
%
\author{Siddharth Pritam}
%
%%%% list of authors for the TOC (use if author list has to be modified)

\institute{Alumni Group of Indian Institute of Technology\\
Kharagpur, India\\
email: sid.iitkgp@gmail.com
}

%\toctitle{Lecture Notes in Computer Science}
%\tocauthor{Authors' Instructions}

\maketitle
\setcounter{footnote}{0}

%
%\begin{abstract}
%
%\end{abstract}
%

\section{Introduction}

 ``Infographics''  is a very new RoboCup Soccer Simulation League 2D team, based on the released source code of agent2d \cite{agent2d} and  Gliders \cite{POWH12,gliders2013tdp}. It also uses librcsc (a base library for The RoboCup Soccer Simulator, RCSS \cite{Manual2002}); soccerwindow2: a viewer program for RCSS; fedit2: a formation editor for agent2d; and GIBBS: in-browser log-player for 2D Simulation League logs \cite{moore}.

The Soccer Simulation 2D League is one of the oldest RoboCup leagues which still presents a significant challenge for Artificial Intelligence, Machine Learning, Autonomous Agents and Multi-Agent Systems, Cognitive Robotics and Complex Systems \cite{Noda2003,balch,IAAI2000,Butler2001,Reis2001,Prokopenko02,Prokopenko03}.  

In this  paper we describe an approach to utilize Fisher information in interaction networks, produced by dynamics of RoboCup Simulation games. This approach is based on (a) methods for computing interaction networks, introduced in \cite{cliff2013towards} and verified  by Gliders \cite{gliders2013tdp}, and (b) methods for determining Fisher information in complex networks, described in \cite{PLOW11}.  It guides optimization of control parameters used by ``Infographics''.

\section{Background}

Information theory \cite{mac03} is an increasingly popular framework for the study of complex systems and their phase transitions \cite{ForemanPW03,liz08c,pro09,Prokopenko13}. In part, this is because complex systems can be viewed as distributed computing systems, and information theory is a natural way to study computation, e.g. \cite{liz08a,liz14a}. Information theory is applicable to any system, provided that one can define probability distribution functions for its states. 

One particular information-theoretic measure, Fisher information, has been extensively used in many fields of science, providing insights that can be directly compared across different system types. It measures the amount of information that an observable random variable has about an unknown parameter, and can also be intuitively interpreted as a measure of the sensitivity of the random variable to changes in the parameter.  For example, a recent study \cite{PLOW11} explicitly related elements of the Fisher information matrix to the rate of change in the
corresponding order parameters, and identified, in particular, second-order phase transitions via divergences of individual elements of the Fisher information matrix.

In parallel, our recent research has also focused on phase transitions between ordered and chaotic behaviour, analyzed from the perspective of distributed computation: specifically, in
small-world Random Boolean Networks (RBN's) \cite{liz11b,liz11b2}. These studies characterized the distributed computation in terms of its
underlying information dynamics: information storage and information transfer, and gave us a good insight on the behavior of the networks and flow of information near critical regimes.

Furthermore, in the context of RoboCup and team sports in general, several new techniques were recently developed for quantifying dynamic interactions in simulated football games \cite{cliff2013towards}. This study not only involved computation of information transfer and storage, but also related the information transfer to responsiveness of the players,
and the information storage within the team to the team's rigidity and lack of tactical flexibility \cite{cliff2013towards}.

We intend to bring these results together, by computing Fisher information within the interaction
networks obtained by the techniques introduced in \cite{cliff2013towards}, and characterising critical behaviour and
phase transitions via Fisher information, following methods of \cite{PLOW11}.
The results promise to advance research in complex multi-agent systems, as well as contribute to
development of new behaviours, implemented within our RoboCup team, ``Infographics'' (India).

\section{Methods}

The first step (1) of the proposed algorithm is computation of ``information-base diagram'', i.e. interaction networks \cite{cliff2013towards}.  It has the following sub-steps, reproduced from \cite{cliff2013towards}:

(1.1) The average conditional transfer entropy \cite{schr00} between any two agents from opposing teams $X$ and $Y$ is computed for a game $g$ with $N$ time steps:
\begin{equation}
    T^g_{Y_i \rightarrow X_j | B}  =  \frac{1}{N} \sum_{n=0}^{N-1} \lim_{k \to \infty} \log_{2}{ \frac{ p(x_{n+1} \mid x^{(k)}_{n},y_{n}, b_n)}{p(x_{n+1} \mid x^{(k)}_{n}, b_n)}} \ ,
\end{equation}
where $Y_i$ is the process tracing the change in the position $y_n$ of agent $i$ ($2 \leq i \leq 11$) from team $Y$ at time $n$ ($1 \leq n \leq N$); $X_j$ is the process tracing the change $x_n$ in the position of  agent $j$ ($2 \leq j \leq 11$) from team $X$ at time $n$; $B$ is the process tracing the change in current ball position $b_n$; and $x^{(k)}_{n} = \{ x_{n-k+1}, \ldots, x_{n-1}, x_n\}$ is the  past of the process $X_j$ (semi-infinite past as $k \to \infty $).

(1.2) The responder agent $J_X(Y_i)$ in team $X$ is identified as the mode of the series 
\[ \{ J_X(Y_i,1), \ldots, J_X(Y_i,G) \} \ , \]
for $G$ games, where
\begin{equation}
  J_X(Y_i,g) =  \argmax_{2 \ \leq \ j \ \leq \ 11}{ \ T^g_{Y_i \rightarrow X_j | B}} \ .
\end{equation}

(1.3) The pairs $(i, J_X)$, where $2 \leq i \leq 11$, identified for each agent $i$ in team $Y$, produce an ``information-base diagram'' $D(Y,X)$ \cite{cliff2013towards}.  There is one outgoing link from each agent $i$ in team $Y$, but there may be several incoming links to the agent $i$ from team $Y$, or an agent $j$ from team $X$.  At this sub-step we select the agent $\hat{J}_X$ in our team $X$ which has the maximal number of incoming links.  In other words, it is the responder agent to the maximal number of opponent agents, being a ``hub'' of the ``information-base diagram''.

\

The next step (2) is computation of Fisher information $F(X_{\hat{J}}|\theta)$ of time series $X_{\hat{J}}$, i.e., the process tracing the change in the hub-agent's position, with respect to some control parameter, $\theta$.  The control parameter may be chosen from a broad set of parameters used within the team.   Varying the control parameter within its range $\Theta$ allows us to compute Fisher information $F(X_{\hat{J}}|\theta)$ for different values of $\theta \in \Theta$, and select the value maximizing Fisher information:
\begin{equation}
\theta^* = \argmax_{\theta \in \Theta} F(X_{\hat{J}}|\theta) \ .
\end{equation}
As argued in \cite{PLOW11}, the value maximizing Fisher information corresponds to a critical point, pinpointing a phase transition.

In summary, this method leads to selection of critical values of control parameters, assisting in optimizing influence of hub-agents on the oppponent team.

\section{Conclusion}

The main motivation behind our approach is the study of phase transitions and relevant control parameters, using interaction networks of RoboCup teams and Fisher information.

We described a method for a selection of control parameters, driven by maximizing Fisher information of a hub-agent in an information-base diagram.  The mechanism is utilized by ``Infographics'', our new simulated soccer team which intends to participate in the RoboCup  Simulation 2D League in 2014.

\subsubsection*{Acknowledgments.}

We thank Hidehisa Akiyama and his colleagues for development of agent2d;  Mikhail Prokopenko, Peter Wang,  and Oliver Obst for their advices regarding the source code of Gliders and helpful discussions of this manuscript; as well as Joseph Lizier for assistance with his Java Information Dynamics toolkit.

\bibliographystyle{splncs}

%\bibliography{refs,references,TE}

\end{document}